OpenStudy (anonymous):

Which situation is best represented by the equation below? 3x = 27 A The total cost of a dinner and tip was $27. If a $3 tip was included in the total cost, what is x, the cost of the dinner? B The total cost of a dinner and dessert was $27. The cost of the dinner was 3 times the cost of the dessert. What is x, the cost of the dessert? C Three friends went to dinner, and each paid $27. What is x, the total amount they spent on dinner? DThree friends went to dinner and paid a total of $27. If they divided the bill evenly, what is x, the amount each friend paid?

OpenStudy (blockcolder):

Do trial and error on all the choices.

OpenStudy (anonymous):

trial and error? whats that

OpenStudy (blockcolder):

It means write out the equation for each choice and see which one matches the given statement.

OpenStudy (anonymous):

oh. for example, what would be the equation for A_?

OpenStudy (anonymous):

@blockcolder

OpenStudy (blockcolder):

For A., the equation is x+3=27.

OpenStudy (anonymous):

because total=add_?

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

@blockcolder because total means add?

OpenStudy (blockcolder):

Yes. :D

OpenStudy (anonymous):

ok.

OpenStudy (campbell_st):

the equation is saying 3 times something is 27 that should help

OpenStudy (anonymous):

let me c if i can figure out B

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

@blockcolder

OpenStudy (anonymous):

I think its B

OpenStudy (anonymous):

because it says the cost eas 3 times the dessert

OpenStudy (blockcolder):

If x is the dessert, then 3x is the dinner and their total is 27. This gives x+3x=27 or 4x=27, so it's not B.

OpenStudy (anonymous):

ur right

OpenStudy (campbell_st):

B says Dinner + dessert = 27 and dinner = 3 x dessert substituting gives 3 dessert + dessert = 37 4 dessert = 27

OpenStudy (anonymous):

it says its D on the answer key, but D says if they divide it uo evenly

OpenStudy (anonymous):

and the equation has to do with multiplying,right?

OpenStudy (blockcolder):

Yeah, so if each friend paid x, we have x+x+x=27, or 3x=27.

OpenStudy (anonymous):

oh.!!!

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

thanks.

OpenStudy (blockcolder):

No problem. :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ill be posting some more questions in a few minutes